OpenStudy (anonymous):

Solve for t. 6t + 3t = 18

pooja195 (pooja195):

Addd 6t+3t=9t 9t=18 divide both side by 9

OpenStudy (anonymous):

t=2

pooja195 (pooja195):

Good job :)

OpenStudy (anonymous):

FIRST: Find two numbers that when multiplied equal -18, and add up to -3 (that is, the coeficients of the second term* and the third term*). Answer = 3 and -6. Because (3)(-6) = -18 and (3)+(-6)=-3 THEN: Multiply two binomials of the form (x + a) (x + b), where "x" will be the square root of the first term* of the original equation, and "a" and "b", will be 3 and -6 (respectively, or the other way around, it doesn't matter) FACTORS: (t + 3) ( t - 6) or (t - 6) (t + 3) PROOF: (t + 3) (t - 6) = t² - 6t + 3t -18 = t² - 3t - 18....or (t - 6) (t + 3) = t² + 3t - 6t -18 = t² - 3t - 18.

OpenStudy (anonymous):

thanks

pooja195 (pooja195):

@Vampirelove28 we arent factoring.

OpenStudy (anonymous):

O ok, Thanks for the heads up